# Correlation between $\mathrm{H}\rm{\alpha}$ emitters and their cosmic web environment at $z\sim 1$

### v1

@@ 5 Conclusions
After performing a few validation tests using dark matter halos and the results from Ramakrishnan et al. (2019) as a benchmark, we apply our approach to mock ELGs extracted from the hydrodynamical suite of simulations IllustrisTNG. For each ELG, we assign the value of $\alpha$ obtained for the parent host halo using the halo mass-dependent smoothing radius advocated by Ramakrishnan et al. (2019). We find that $>90$% of the luminous ELGs with $L_{\mathrm{H}\rm{\alpha}}\geq 10^{42}\ {\rm erg\,s^{-1}}$ reside in knots (low $\alpha$) and filaments (intermediate $\alpha$), while the rest occupies sheets (high $\alpha$). ELGs with a larger $\alpha$ are more strongly biased on large scales (i.e. their $b_{1}$ is larger), as expected.
@@
We have quantified how galaxy properties — including the different physical sources of $\mathrm{H}\rm{\alpha}$ emission and the $\mathrm{H}\rm{\alpha}$ line FWHM — correlate with $\alpha$ using Spearman’s rank correlation coefficient $\rho_{s}$. We split our results between central and satellite galaxies. All the correlations are weak ($|\rho_{s}|<0.5$), but they are statistically significant up to a host halo mass $M=10^{12}\ M_{\odot}/h$ when measured as a function of $M$. They provide clues for the physics of galaxy formation, at least, as it is implemented in IllustrisTNG. At fixed halo mass, the correlation between RR-induced $\mathrm{H}\rm{\alpha}$ emission from central galaxies and $\alpha$ is the strongest for $M\lesssim 10^{12}\ M_{\odot}/h$. This correlation, which vanishes nearly entirely when centrals are selected by total $\mathrm{H}\rm{\alpha}$ luminosity, is presumably related to the formation of cosmological-scale accretion shocks in sheets and filaments (Mo et al., 2010; Mandelker et al., 2019; Ramsøy et al., 2021; Pasha et al., 2023; Hasan et al., 2023, 2024). For satellite galaxies, $\alpha$ exhibits the strongest correlation with $L_{\mathrm{H}\rm{\alpha}}^{\text{AGN}}$. The correlations between galaxy properties and anisotropy parameter are also reflected in the scale-dependence of the ELG bias $b(k)$. For central galaxies with $\mathrm{H}\rm{\alpha}$ luminosity $L_{\mathrm{H}\rm{\alpha}}\geq 10^{41}\ {\rm erg\,s^{-1}}$ for example, the scale-dependence of $b(k)$ is weaker for those with high $L_{\mathrm{H}\rm{\alpha}}$, SFR or $\mathrm{H}\rm{\alpha}$ FWHM.
@@
Hadzhiyska et al. (2021) performed a similar analysis, albeit using standard luminosity-SFR relations (e.g., Kennicutt, 1983) to compute the $\mathrm{H}\rm{\alpha}$ luminosity of mock galaxies in IllustrisTNG-300-1. More specifically, they considered the specific star formation rate (sSFR) given by the ratio of SFR to total stellar mass Furthermore, they adopted a fixed filtering scale of $4\ {\rm{\it h}^{-1}Mpc}$ significantly larger than our mass-dependent $\text{R}_{\text{G}}(M)$ — which varies in the range $0.25\lesssim\rm{R_{G}}\lesssim 2\ {\rm{\it h}^{-1}Mpc}$ for halos in mass range $2\times 10^{11}<M<10^{14}\ {\rm{\it h}}^{-1}M_{\odot}$ — to classify galaxies into knots, filaments, etc. Finally, they did not work with the anisotropy parameter $\alpha$ to quantify the correlation between large scale clustering and local environment, but considered instead more traditional proxies such as the environment density. Nonetheless, our conclusions are consistent with theirs wherever there is overlap.
@@
Would it be possible to select ELGs with $\mathrm{H}\rm{\alpha}$ emission dominated by a specific source (e.g., RR, AGN, etc.) using other spectral signatures? For $\mathrm{H}\rm{\alpha}$ emitters dominated by AGN-driven emission, one would also expect significant $\mathrm{H}\rm{\alpha}$ emission from what is referred to as their broad (and narrow) line region BLR (NLR), which our model does not fully take into account. Rapoport et al. (2025) showed the minor effect of excluding AGN (a small fraction) from the calculation. In reality, bright AGNs are obvious, but low luminosity or obscured AGN where the broad lines are not detected unambiguously are harder to distinguish from non-AGN galaxies, and one can think of excluding the central pixel of the galaxy to exclude AGN contributions.
@@
Finally, the strong dependence of $b(k)$ on the environment could prove useful for multi-tracer constraints on cosmology (e.g., Seljak, 2009; McDonald & Seljak, 2009; White et al., 2009; Hamaus et al., 2010, 2011; Cai et al., 2011; Blake et al., 2013; Abramo & Leonard, 2013; Ferramacho et al., 2014; Abramo et al., 2016; Yamauchi et al., 2017; Zhao et al., 2021; Schaan & White, 2021; Sullivan et al., 2023; Heinrich et al., 2024; Montero-Dorta & Rodriguez, 2024; Euclid Collaboration et al., 2024c; Fang et al., 2024). To maximize the gains of this approach, the product $(\bar{n}P)$ should (vastly) exceed unity for each tracer. Using both the $\alpha_{\rm V}$ and $\alpha_{\rm T}$ classifications, we find $\bar{n}P\gtrsim 10$ (resp. $\bar{n}P\gtrsim 100$) at $k\sim k_{\text{eq}}$ for $z=1$ ELGs with luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$ (resp. $L_{\mathrm{H}\rm{\alpha}}>10^{41}\ {\rm erg\,s^{-1}}$) residing in knots or filaments.
@@
Multi-tracer analyses are expected to be particularly useful for constraining the level of primordial non-Gaussianity in the initial conditions (Hamaus et al., 2011). Previous work on this problem has looked at environment defined by large-scale density only, finding only modest gains in constraining power (Morawetz et al., 2025). This is thought to be because the tracers populate only a narrow range of halo masses (the most massive halos). Our results provide strong motivation for using $\mathrm{H}\rm{\alpha}$ emitters to constrain primordial non-Gaussianity because we have shown that (a) $\mathrm{H}\rm{\alpha}$ emitters span a wide range of halo masses and (b) they present a large environmental dependence that is not confined to density alone.
6 Acknowledgements
@@
We thank Shmuel Bialy, Nir Mandelker, Adi Nusser, Aseem Paranjape and Sujatha Ramakrishnan for helpful discussions and comments on the manuscript. I.R, V.D. and E.B. acknowledge support from the Israel Science Foundation (grant no. 2562/20).
@@ Appendix A Full Width Half Maximum
where $L_{\mathrm{H}\rm{\alpha}}^{\text{i}}$ is the $\mathrm{H}\rm{\alpha}$ luminosity in the cell. The intrinsic spectrum of a galaxy is the sum of all contributing cells $L_{\lambda,\mathrm{H}\rm{\alpha}}=\sum_{\text{i}}L_{\lambda,\mathrm{H}\rm{\alpha}}^{\text{i}}$. The sampling wavelengths $\lambda$ are constructed in an adaptive manner such that no contributions are missed, and the sampling resolution is fixed to $1\rm{\AA}$. Finally, the intrinsic spectrum is smoothed using a Gaussian filter with a spectral resolution $\lambda/\Delta\lambda=400$ matching the resolving power of the Euclid NISP spectrometer (Euclid Collaboration et al., 2024a, d). Due to this moderate spectral resolution, the resulting smoothed line profiles are mostly featureless except for a single peak with a well-defined FWHM.
@@ Appendix B Numerical convergence
Figure 13: Spearman correlation coefficient $\rho_{s}$ for the correlation between $\alpha_{\rm{V}}$ and the RR-, AGN-induced $\mathrm{H}\rm{\alpha}$ luminosities as well as the $\mathrm{H}\rm{\alpha}$ line FWHM for central galaxies. Results are shown for TNG50-1 (solid curves) and TNG300-1 (dashed curves).

### v2

@@ V Conclusions
After performing a few validation tests using dark matter halos and the results from Ramakrishnan et al. (2019) as a benchmark, we apply our approach to mock ELGs extracted from the hydrodynamical suite of simulations IllustrisTNG. For each ELG, we assign the value of $\alpha$ obtained for the parent host halo using the halo mass-dependent smoothing radius advocated by Ramakrishnan et al. (2019). We find that >90>90% of the luminous ELGs with $L_{\mathrm{H}\rm{\alpha}}\geq 10^{42}\ {\rm erg\,s^{-1}}$ reside in knots (low $\alpha$) and filaments (intermediate $\alpha$), while the rest occupies sheets (high $\alpha$). ELGs with a larger $\alpha$ are more strongly biased on large scales (i.e. their $b_{1}$ is larger), as expected.
@@
We have quantified how galaxy properties — including the different physical sources of $\mathrm{H}\rm{\alpha}$ emission and the $\mathrm{H}\rm{\alpha}$ line FWHM — correlate with $\alpha$ using Spearman’s rank correlation coefficient $\rho_{s}$. We split our results between central and satellite galaxies. All the correlations are weak ($|\rho_{s}|<0.5$), but they are statistically significant up to a host halo mass $M=10^{12}\ M_{\odot}/h$ when measured as a function of $M$. They provide clues for the physics of galaxy formation, at least, as it is implemented in IllustrisTNG. At fixed halo mass, the correlation between RR-induced $\mathrm{H}\rm{\alpha}$ emission from central galaxies and $\alpha$ is the strongest for $M\lesssim 10^{12}\ M_{\odot}/h$. This correlation, which vanishes nearly entirely when centrals are selected by total $\mathrm{H}\rm{\alpha}$ luminosity, is related to the formation of cosmological-scale accretion shocks in sheets and filaments (Mo et al., 2010; Mandelker et al., 2019; Ramsøy et al., 2021; Pasha et al., 2023; Hasan et al., 2023, 2024). For satellite galaxies, $\alpha$ exhibits the strongest correlation with $L_{\mathrm{H}\rm{\alpha}}^{\text{AGN}}$. The correlations between galaxy properties and anisotropy parameter are also reflected in the scale-dependence of the ELG bias $b(k)$. For central galaxies with $\mathrm{H}\rm{\alpha}$ luminosity $L_{\mathrm{H}\rm{\alpha}}\geq 10^{41}\ {\rm erg\,s^{-1}}$ for example, the scale-dependence of $b(k)$ is weaker for those with high $L_{\mathrm{H}\rm{\alpha}}$, SFR or $\mathrm{H}\rm{\alpha}$ FWHM.
@@
Hadzhiyska et al. (2021) performed a similar analysis, albeit using standard luminosity-SFR relations (e.g., Kennicutt, 1983) to compute the $\mathrm{H}\rm{\alpha}$ luminosity of mock galaxies in IllustrisTNG-300-1. More specifically, they considered the specific star formation rate (sSFR) given by the ratio of SFR to total stellar mass
Furthermore, they adopted a fixed filtering scale of $4\ {\rm{\it h}^{-1}Mpc}$ significantly larger than our mass-dependent $\text{R}_{\text{G}}(M)$ — which varies in the range $0.25\lesssim\rm{R_{G}}\lesssim 2\ {\rm{\it h}^{-1}Mpc}$ for halos in mass range $2\times 10^{11}<M<10^{14}\ {\rm{\it h}}^{-1}M_{\odot}$ — to classify galaxies into knots, filaments, etc. Finally, they did not work with the anisotropy parameter $\alpha$ to quantify the correlation between large scale clustering and local environment, but considered instead more traditional proxies such as the environment density. Nonetheless, our conclusions are consistent with theirs wherever there is overlap.
@@
Would it be possible to select ELGs with $\mathrm{H}\rm{\alpha}$ emission dominated by a specific source (e.g., RR, AGN, etc.) using other spectral signatures? For ELGs dominated by AGN-driven emission, one would also expect significant $\mathrm{H}\rm{\alpha}$ emission from what is referred to as their broad (and narrow) line region BLR (NLR), which our model does not fully take into account. Rapoport et al. (2025) showed the minor effect of excluding AGN (a small fraction) from the calculation. In reality, bright AGNs are obvious, but low luminosity or obscured AGN where the broad lines are not detected unambiguously are harder to distinguish from non-AGN galaxies, and one can think of excluding the central pixel of the galaxy to exclude AGN contributions.
@@
Finally, the strong dependence of $b(k)$ on the environment could prove useful for multi-tracer constraints on cosmology (e.g., Seljak, 2009; McDonald and Seljak, 2009; White et al., 2009; Hamaus et al., 2010, 2011; Cai et al., 2011; Blake et al., 2013; Abramo and Leonard, 2013; Ferramacho et al., 2014; Abramo et al., 2016; Yamauchi et al., 2017; Zhao et al., 2021; Schaan and White, 2021; Sullivan et al., 2023; Heinrich et al., 2024; Montero-Dorta and Rodriguez, 2024; Euclid Collaboration et al., 2024b; Fang et al., 2024). To maximize the gains of this approach, the product $(\bar{n}P)$ should (vastly) exceed unity for each tracer. Using both the $\alpha_{\rm V}$ and $\alpha_{\rm T}$ classifications, we find $\bar{n}P\gtrsim 10$ (resp. $\bar{n}P\gtrsim 100$) at $k\sim k_{\text{eq}}$ for $z=1$ ELGs with luminosity $L_{\mathrm{H}\rm{\alpha}}>10^{42}\ {\rm erg\,s^{-1}}$ (resp. $L_{\mathrm{H}\rm{\alpha}}>10^{41}\ {\rm erg\,s^{-1}}$) residing in knots or filaments.
@@
Multi-tracer analyses are expected to be particularly useful for constraining the level of primordial non-Gaussianity in the initial conditions (Hamaus et al., 2011). Previous work on this problem has looked at environment defined by large-scale density only, finding only modest gains in constraining power (Morawetz et al., 2025). This is thought to be because the tracers populate only a narrow range of halo masses (the most massive halos). Our results provide strong motivation for using ELGs to constrain primordial non-Gaussianity because we have shown that (a) $\mathrm{H}\rm{\alpha}$ emitters span a wide range of halo masses and (b) they present a large environmental dependence that is not confined to density alone.
VI Acknowledgements
@@
We thank Shmuel Bialy, Nir Mandelker, Adi Nusser, Aseem Paranjape and Sujatha Ramakrishnan for helpful discussions and comments on the manuscript. The Technion team acknowledges support from the Israel Science Foundation (grant nos. 2562/20 and 2617/25).
@@ Appendix A Full Width Half Maximum
where $L_{\mathrm{H}\rm{\alpha}}^{\text{i}}$ is the $\mathrm{H}\rm{\alpha}$ luminosity in the cell. The intrinsic spectrum of a galaxy is the sum of all contributing cells $L_{\lambda,\mathrm{H}\rm{\alpha}}=\sum_{\text{i}}L_{\lambda,\mathrm{H}\rm{\alpha}}^{\text{i}}$. The sampling wavelengths $\lambda$ are constructed in an adaptive manner such that no contributions are missed, and the sampling resolution is fixed to $1\rm{\AA }$. Finally, the intrinsic spectrum is smoothed using a Gaussian filter with a spectral resolution $\lambda/\Delta\lambda=400$ matching the resolving power of the Euclid NISP spectrometer (Euclid Collaboration et al., 2024d, c). Due to this moderate spectral resolution, the resulting smoothed line profiles are mostly featureless except for a single peak with a well-defined FWHM.
@@ Appendix B Numerical convergence
Figure 13: Spearman correlation coefficient $\rho_{s}$ for the correlation between $\alpha_{\rm{V}}$ and the $R$-, AGN-induced $\mathrm{H}\rm{\alpha}$ luminosities as well as the $\mathrm{H}\rm{\alpha}$ line FWHM for central galaxies. Results are shown for TNG50-1 (solid curves) and TNG300-1 (dashed curves).